# Collapse or Thrive? Perils and Promises of Synthetic Data in a Self-Generating World

### v1

@@ 6 Future Directions
Appendix A Iterative Gaussian Model Fitting: Mathematical Results and Proofs
A.1 Setup
@@
Appendix B Iterative KDE Fitting: Mathematical Results and Proofs
@@
Suppose that data accumulates as in Section 2.2 for a Gaussian kernel. Let the bandwidth at the $n$th model-fitting iteration be $c(tn)^{-1/5}$ for a constant $c$. Then the asymptotic variance of the limiting KDE is finite.
@@
Appendix C Experimental Results: Sweep Configurations
C.1 Model Collapse in Multivariate Gaussian Modeling
@@
"Accumulate",
"Accumulate-Subsample",
"Replace",
@@
C.2 Model Collapse in Kernel Density Estimation
Seeds were swept from 0 to 99, inclusive.
C.3 Model Collapse in Kernel Density Estimation
@@
"Accumulate",
"Accumulate-Subsample",
"Replace",
@@
"Accumulate",
"Accumulate-Subsample",
"Replace",
@@
"Accumulate",
"Accumulate-Subsample",
"Replace",
@@
values: ["swiss_roll"]
@@
"Accumulate",
"Accumulate-Subsample",
"Replace",
@@
C.4 Model Collapse in Linear Regression
@@
"Accumulate",
"Accumulate-Subsample",
"Replace",
@@
Appendix D Additional Experimental Results for Model Collapse Hyperparameters
@@
D.1 Additional Results for Model Collapse in Linear Regression

### v4

@@ Appendix A Related Work
Appendix B Iterative Gaussian Model Fitting: Mathematical Results and Proofs
B.1 Setup
@@
Appendix C Iterative KDE Fitting: Mathematical Results and Proofs
@@
Under the accumulate training workflow consider density estimation as in Section 2.2 with a Gaussian kernel. Let the bandwidth at the $t$th model-fitting iteration be $c\cdot(tn)^{-1/5}$ for a constant $c$. Then the asymptotic variance of the limiting KDE is finite.
@@
Appendix D Experimental Results: Sweep Configurations
D.1 Model Collapse in Multivariate Gaussian Modeling
@@
Accumulate”,
Accumulate-Subsample”,
Replace”,
@@
D.2 Model Collapse in Kernel Density Estimation
@@
Accumulate”,
Accumulate-Subsample”,
Replace”,
@@
Accumulate”,
Accumulate-Subsample”,
Replace”,
@@
Accumulate”,
Accumulate-Subsample”,
Replace”,
@@
values: [”swiss_roll”]
@@
Accumulate”,
Accumulate-Subsample”,
Replace”,
@@
D.3 Model Collapse in Linear Regression
@@
Accumulate”,
Accumulate-Subsample”,
Replace”,
@@
Appendix E Additional Experimental Results for Model Collapse Hyperparameters
@@
E.1 Additional Results for Model Collapse in Linear Regression